OpenStudy (purnanies):

Radical Expressions : 5 square root of 6 / 2 square root of 75.... Answer is square root of 2 divided by 2, but I need a step by step explanation

OpenStudy (3mar):

Need help?

OpenStudy (purnanies):

yis

OpenStudy (peachpi):

\[\frac{ 5\sqrt{6} }{ 2\sqrt{75} }=\frac{ 5\sqrt{2\times3} }{ 2\sqrt{25\times3} }=\frac{ 5\sqrt{2}\sqrt{3} }{ 2\sqrt{25}\sqrt{3} }=\frac{5\sqrt{2}\sqrt{3}}{2\times5\sqrt3}=\frac{\sqrt{2}}{2}\]

OpenStudy (3mar):

peachpi did it very well!